# A globalized inexact semismooth Newton method for strongly convex optimal control problems

### v3

@@ 5 Semismooth Newton
Let $(\xi_{k})$ be a sequence of iterates of Algorithm 2. Then it holds $\liminf_{k\to\infty}t_{k}>0$ and $\nabla\Phi(\xi_{k})\to 0$ in $Y$.
Proof.
By Lemma 5.7, we have $t_{k}\langle d_{k},\nabla\Phi(\xi_{k})\rangle\to 0$. If $\liminf_{k\to\infty}t_{k}>0$ then (5.7) implies $\|\nabla\Phi(\xi_{k})\|_{Y}\to 0$. Consider the case $\liminf_{k\to\infty}t_{k}=0$. By extracting a subsequence, we can assume $t_{k}\to 0$ and $t_{k}<1$ for all $k$. Then from $t_{k}\to 0$ and $t_{k}\langle d_{k},\nabla\Phi(\xi_{k})\rangle\to 0$ we get with the help of Lemma 5.8
@@
Dividing by $\|d_{k}\|_{Y}$ and passing to the limit $k\to\infty$ leads to the contradiction $1\leq 0$. Here, we used $\tau\leq 1$, and $\frac{a^{1+\tau}}{b}=\left(\frac{a^{2}}{b}\right)^{\tau}\left(\frac{a}{b}\right)^{1-\tau}$. ∎
@@
Let $(\xi_{k})$ be a sequence of iterates of Algorithm 2. Then $\xi_{k}\to\bar{\xi}$ in $Y$, where $\xi$ is the unique solution of $\Phi(\xi)=0$. In addition, $d_{k}\to 0$ in $Y$.
Proof.
Let $\bar{\xi}$ be such that $\Phi(\bar{\xi})=0$. The existence of such $\bar{\xi}$ follows from Lemma 2.7. Lemma 5.5 implies $\|\bar{\xi}-\xi_{k}\|_{Y}\leq\|\nabla\Phi(\xi_{k})\|_{Y}$, the latter quantity tends to zero by Lemma 5.9. ∎
The next step is to prove that the step size $t_{k}=1$ is accepted for all $k\geq k_{0}$. Here, we follow the exposition in [7]. First, we establish superlinear convergence of the local inexact semismooth Newton method.
@@
Let $\bar{\xi}$ be the unique solution of $\Phi(\bar{\xi})=0$. Let $(\xi_{k})$ be a sequence of iterates of Algorithm 2 with step sizes $(t_{k})$. Let $r_{k}:=M_{k}d_{k}+\nabla\Phi(\xi_{k})$. Then
@@
Proof.
Using the inexact Newton equation (5.5) and Corollary 5.1, we get
@@
where we have used Corollary 5.3 and the semismoothness of $\nabla\Phi$. ∎
Now, we will prove that the step size $t_{k}=1$ will be accepted for large $k$. This proof crucially relies on the second-order expansion in the definition of second-order semismoothness (Definition 3.1).
@@
Let $\bar{\xi}$ be the unique solution of $\Phi(\bar{\xi})=0$. Let $(\xi_{k})$ be a sequence of iterates of Algorithm 2 with step sizes $(t_{k})$. Then there is $k_{0}>0$ such that $t_{k}=1$ for all $k\geq k_{0}$.
Proof.
See also [7, Theorem 3.3]. Using the simple inequalities
@@
$\Phi(\xi_{k}+d_{k})-\Phi(\bar{\xi})-\langle\nabla\Phi(\bar{\xi}),\xi_{k}+d_{k}-\bar{\xi}\rangle\\ \begin{aligned} &=-\frac{1}{2}\langle H_{k}(\xi_{k}+d_{k}-\bar{\xi}),\xi_{k}+d_{k}-\bar{\xi}\rangle+o(\|\xi_{k}+d_{k}-\bar{\xi}\|_{Y}^{2})\\ &=o(\|\xi_{k}-\bar{\xi}\|_{Y}^{2}),\end{aligned}$
where $H_{k}\in\partial^{2}\Phi(\bar{\xi}_{k}+d_{k})$, which is bounded due to Corollary 5.2. Similarly, we prove using $\nabla\Phi(\bar{\xi})=0$
@@
$\Phi(\xi_{k}+d_{k})-\Phi(\xi_{k})-\sigma\langle d_{k},\nabla\Phi(\xi_{k})\rangle\\ =\left(\frac{1}{2}-\sigma\right)\langle d_{k},\nabla\Phi(\xi_{k})\rangle+o(\|\xi_{k}-\bar{\xi}\|_{Y}^{2})\\ =-\left(\frac{1}{2}-\sigma\right)\|\xi_{k}-\bar{\xi}\|_{Y}^{2}+o(\|\xi_{k}-\bar{\xi}\|_{Y}^{2}).$
The right-hand side is less than zero for $k$ large enough, which proves the claim. ∎
@@
Let 1 be satisfied. Let $\bar{\xi}$ be the unique solution of $\nabla\Phi(\bar{\xi})=0$. Let $(\xi_{k})$ be a sequence of iterates of Algorithm 2. Then $\xi_{k}\to\bar{\xi}$ q-superlinearly.
Proof.
Due to Lemma 5.12, there is $k_{0}>0$ such that $t_{k}=1$ for all $k\geq k_{0}$. Then Lemma 5.11 implies $\|\xi_{k}+d_{k}-\bar{\xi}\|_{Y}=\|\xi_{k+1}-\bar{\xi}\|_{Y}=o(\|\xi_{k}-\bar{\xi}\|_{Y})\|$ for $k\geq k_{0}$, which is the claim. ∎
@@
The proof of transition to fast convergence in [18, Lemma 4.4] relied on stronger assumptions. There a condition on the measure of almost active sets was used, which implies strict complementarity. In addition, the operator $S^{*}$ has to be continuous with values in $C^{1}(\bar{\Omega})$. Under these conditions $\partial^{2}\Psi$ is single-valued and Hölder continuous in a neighborhood of the solution [18, Lemma 3.4], which allows to prove the higher-order convergence rate $\|\xi_{k+1}-\bar{\xi}\|_{Y}\leq c\|\xi_{k}-\bar{\xi}\|_{Y}^{3/2}$.
@@
$0=\bar{u}-\operatorname{prox}_{g/\alpha}(-S^{*}(S\bar{u}-z)/\alpha).$ (5.8)
@@
$\delta u+\frac{1}{\alpha}H_{k}S^{*}S\delta u=-\left(u_{k}-\operatorname{prox}_{g/\alpha}(-S^{*}(Su_{k}-z)/\alpha)\right).$ (5.9)
@@
The map $p\mapsto p-S^{*}(S\operatorname{prox}_{g/\alpha}(-p/\alpha)-z)$ is called normal map after [33]. Semismooth Newton methods based on this equation are considered in [27, 17]. In the recent article [26], the convergence analysis of a trust-region globalization of a semismooth Newton method applied to the normal map equation is developed in finite dimensions, where globalization is based on the residual in the normal equation.
@@
In [10] a globalization of the active-set strategy [3] to solve discretized box constrained optimal control problems was proposed. However, the analysis does not carry over to the infinite-dimensional setting, as certain operators are not invertible in infinite dimensions [10, Remark 3].
@@
where $F:Y\to\mathbb{R}$ is convex. Let $F^{*}$ denote the convex conjugate of $F$. Under the assumption that $F$ is strongly convex and twice continuously differentiable, we get that $F^{*}$ is also strongly convex and twice continuously differentiable. Let us denote with $\nabla F^{*}:=\nabla(F^{*})$ its gradient. Then, we have $\nabla F(\nabla F^{*}(y))=y$ for all $y\in Y$ by [2, Theorem 16.29]. Differentiating this identity yields $\nabla^{2}F(\nabla F^{*}(y))\nabla^{2}F^{*}(y))=\operatorname{id}$ or $\nabla^{2}F^{*}(y)=\nabla^{2}F(\nabla F^{*}(y))^{-1}$. The resulting second-order derivative of $\Phi$ is given by $\partial^{2}\Phi=\nabla^{2}F^{*}+\partial^{2}\Psi$. Then all the convergence analysis for the corresponding semismooth Newton algorithm remains true, only the constants in Lemma 5.5 and Corollary 5.3 have to be adapted.
@@ 6 Semismoothness of Nemyzki operators
$g(u)=\int_{\Omega}\tilde{g}(u(x))\,\text{{}{}d}\mu$ (6.1)
@@
Proof.
This follows from the results of [34, Section 3], in particular the characterization of $\partial g(u)$ as
@@ 6.1 Piecewise smooth proximals
Let $\operatorname{prox}_{\tilde{g}/\alpha}$ satisfy 2. Define the single-valued map
@@
Proof.
Take $v\in\mathbb{R}$, and $h$ such that $\operatorname{prox}_{\tilde{g}/\alpha}$ is continuously differentiable between $v$ and $v+h$. Then both claims follow easily. ∎
Clearly, this result can be generalized for $\operatorname{prox}_{\tilde{g}/\alpha}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with infinitely many pieces, see also [40, Section 2.5.3].
In the next step, we prove that the derivative of the proximal operator of the integral functional $g$ is semismooth from $L^{q}(\Omega)$ to $L^{p}(\Omega)$ for $p<q$. The proof of the following result is adapted from the proof of [30, Proposition 6], more general results can be found in [40, Section 3.3].
@@
Let $1\leq p<q\leq\infty$ be given. Let $\mu(\Omega)<\infty$. Let $\operatorname{prox}_{\tilde{g}/\alpha}$ satisfy 2. Let $v\in L^{q}(\Omega)$. Define the measurable function $\partial\operatorname{prox}_{g/\alpha}(v)$ for $\mu$-almost all $x\in\Omega$ by
@@
In addition, we set $r(x,0)=0$ for $x\in\Omega$. Then $x\mapsto r(x,h(x))$ is $\mu$-measurable for all measurable $h:\Omega\to\mathbb{R}$. In addition, $\lim_{h\to 0}r(x,h)=0$ for almost all $x$ by Lemma 6.2. Since $\operatorname{prox}_{\tilde{g}/\alpha}$ is Lipschitz (Proposition 2.4) and consequently $|\partial\operatorname{prox}_{\tilde{g}/\alpha}|\leq 1$, we have for almost all $x$ and all $h$ the estimate $|r(x,h)|\leq 2$.
@@
The second claim can be proven with similar arguments, observing that $|\operatorname{prox}_{\tilde{g}/\alpha}^{\prime}(v;h)|\leq|h|$. ∎
Note that we cannot get rid of the norm gap in the previous result. The necessity of this gap has been discussed in [40, Example 3.57], [19, Example 8.14] for semismoothness and in [11, Remark 4.4], [24] for Bouligand differentiability.
@@
Without the assumption $\mu(\Omega)<\infty$, the result of Lemma 6.3 is not true. Let
@@
Due to the norm gap appearing in Lemma 6.3, we cannot expect that $\operatorname{prox}_{g/\alpha}$ is semismooth as a mapping on $L^{2}(\Omega)$. In order that 1 is satisfied, we will need to rely on smoothing properties of $S$ or $S^{*}$.
@@
Assume $\mu(\Omega)<\infty$. Set $U:=L^{2}(\Omega)$. Let $\operatorname{prox}_{\tilde{g}/\alpha}$ satisfy 2. Let $Y$ be a Hilbert space, $S\in\mathcal{L}(L^{2}(\Omega),Y)$ be given such that the Hilbert space adjoint $S^{*}$ belongs to $\mathcal{L}(Y,L^{q}(\Omega))$ for some $q>2$.
@@
$\partial^{2}\Psi(\xi):=\left\{\frac{1}{\alpha}S\partial\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)S^{*}\right\},$ (6.2)
where $\partial\operatorname{prox}_{g/\alpha}$ is defined in Lemma 6.3. Moreover, 1 is satisfied.
@@
Proof of Lemma 6.5.
We show that $\partial^{2}\Psi(\xi)$ from (6.2) satisfies the assumptions of Lemma 3.2. Let us recall $\nabla\Psi(\xi)=S\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)$. Let $v,h\in Y$ be given. Then Lemma 6.3 with $p=2$ and the assumption on $S^{*}$ imply
@@
Clearly this implies that $\partial^{2}\Psi(\xi)$ satisfies (3.2). The Bouligand property of the directional derivative can be proven with analogous arguments. ∎

### v4

@@ 5 Semismooth Newton
Let $(\xi_{k})$ be a sequence of iterates of Algorithm˜2. Then it holds $\liminf_{k\to\infty}t_{k}>0$ and $\nabla\Phi(\xi_{k})\to 0$ in $Y$.
Proof 5.16.
By Lemma˜5.11, we have $t_{k}\langle d_{k},\nabla\Phi(\xi_{k})\rangle\to 0$. If $\liminf_{k\to\infty}t_{k}>0$ then (5.7) implies $\|\nabla\Phi(\xi_{k})\|_{Y}\to 0$. Consider the case $\liminf_{k\to\infty}t_{k}=0$. By extracting a subsequence, we can assume $t_{k}\to 0$ and $t_{k}<1$ for all $k$. Then from $t_{k}\to 0$ and $t_{k}\langle d_{k},\nabla\Phi(\xi_{k})\rangle\to 0$ we get with the help of Lemma˜5.13
@@
Dividing by $\|d_{k}\|_{Y}$ and passing to the limit $k\to\infty$ leads to the contradiction $1\leq 0$. Here, we used $\tau\leq 1$, and $\frac{a^{1+\tau}}{b}=\left(\frac{a^{2}}{b}\right)^{\tau}\left(\frac{a}{b}\right)^{1-\tau}$.
@@
Let $(\xi_{k})$ be a sequence of iterates of Algorithm˜2. Then $\xi_{k}\to\bar{\xi}$ in $Y$, where $\xi$ is the unique solution of $\Phi(\xi)=0$. In addition, $d_{k}\to 0$ in $Y$.
Proof 5.18.
Let $\bar{\xi}$ be such that $\Phi(\bar{\xi})=0$. The existence of such $\bar{\xi}$ follows from Lemma˜2.9. Lemma˜5.7 implies $\|\bar{\xi}-\xi_{k}\|_{Y}\leq\|\nabla\Phi(\xi_{k})\|_{Y}$, the latter quantity tends to zero by Lemma˜5.15.
The next step is to prove that the step size $t_{k}=1$ is accepted for all $k\geq k_{0}$. Here, we follow the exposition in [Facchinei1995]. First, we establish superlinear convergence of the local inexact semismooth Newton method.
@@
Let $\bar{\xi}$ be the unique solution of $\Phi(\bar{\xi})=0$. Let $(\xi_{k})$ be a sequence of iterates of Algorithm˜2 with step sizes $(t_{k})$. Let $r_{k}:=M_{k}d_{k}+\nabla\Phi(\xi_{k})$. Then
@@
Proof 5.20.
Using the inexact Newton equation (5.5) and Corollary˜5.1, we get
@@
where we have used Corollary˜5.4 and the semismoothness of $\nabla\Phi$.
Now, we will prove that the step size $t_{k}=1$ will be accepted for large $k$. This proof crucially relies on the second-order expansion in the definition of second-order semismoothness (Definition˜3.1).
@@
Let $\bar{\xi}$ be the unique solution of $\Phi(\bar{\xi})=0$. Let $(\xi_{k})$ be a sequence of iterates of Algorithm˜2 with step sizes $(t_{k})$. Then there is $k_{0}>0$ such that $t_{k}=1$ for all $k\geq k_{0}$.
Proof 5.22.
See also [Facchinei1995, Theorem 3.3]. Using the simple inequalities
@@
$\Phi(\xi_{k}+d_{k})-\Phi(\bar{\xi})-\langle\nabla\Phi(\bar{\xi}),\xi_{k}+d_{k}-\bar{\xi}\rangle$ $=-\frac{1}{2}\langle H_{k}(\xi_{k}+d_{k}-\bar{\xi}),\xi_{k}+d_{k}-\bar{\xi}\rangle+o(\|\xi_{k}+d_{k}-\bar{\xi}\|_{Y}^{2})$
$=o(\|\xi_{k}-\bar{\xi}\|_{Y}^{2}),$
where $H_{k}\in\partial^{2}\Phi(\bar{\xi}_{k}+d_{k})$, which is bounded due to Corollary˜5.3. Similarly, we prove using $\nabla\Phi(\bar{\xi})=0$
@@
$\Phi(\xi_{k}+d_{k})-\Phi(\xi_{k})-\sigma\langle d_{k},\nabla\Phi(\xi_{k})\rangle$ $=\left(\frac{1}{2}-\sigma\right)\langle d_{k},\nabla\Phi(\xi_{k})\rangle+o(\|\xi_{k}-\bar{\xi}\|_{Y}^{2})$
$=-\left(\frac{1}{2}-\sigma\right)\|\xi_{k}-\bar{\xi}\|_{Y}^{2}+o(\|\xi_{k}-\bar{\xi}\|_{Y}^{2}).$
The right-hand side is less than zero for $k$ large enough, which proves the claim.
@@
Let ˜1 be satisfied. Let $\bar{\xi}$ be the unique solution of $\nabla\Phi(\bar{\xi})=0$. Let $(\xi_{k})$ be a sequence of iterates of Algorithm˜2. Then $\xi_{k}\to\bar{\xi}$ q-superlinearly.
Proof 5.24.
Due to Lemma˜5.21, there is $k_{0}>0$ such that $t_{k}=1$ for all $k\geq k_{0}$. Then Lemma˜5.19 implies $\|\xi_{k}+d_{k}-\bar{\xi}\|_{Y}=\|\xi_{k+1}-\bar{\xi}\|_{Y}=o(\|\xi_{k}-\bar{\xi}\|_{Y})\|$ for $k\geq k_{0}$, which is the claim.
@@
The proof of transition to fast convergence in [HinzeVierling2012oms, Lemma 4.4] relied on stronger assumptions. There a condition on the measure of almost active sets was used, which implies strict complementarity. In addition, the operator $S^{*}$ has to be continuous with values in $C^{1}(\bar{\Omega})$. Under these conditions $\partial^{2}\Psi$ is single-valued and Hölder continuous in a neighborhood of the solution [HinzeVierling2012oms, Lemma 3.4], which allows to prove the higher-order convergence rate $\|\xi_{k+1}-\bar{\xi}\|_{Y}\leq c\|\xi_{k}-\bar{\xi}\|_{Y}^{3/2}$.
@@
(5.8) $0=\bar{u}-\operatorname{prox}_{g/\alpha}(-S^{*}(S\bar{u}-z)/\alpha).$
@@
(5.9) $\delta u+\frac{1}{\alpha}H_{k}S^{*}S\delta u=-\left(u_{k}-\operatorname{prox}_{g/\alpha}(-S^{*}(Su_{k}-z)/\alpha)\right).$
@@
The map $p\mapsto p-S^{*}(S\operatorname{prox}_{g/\alpha}(-p/\alpha)-z)$ is called normal map after [Robinson1992]. Semismooth Newton methods based on this equation are considered in [PieperDiss, HinzeVierling_2012spp]. In the recent article [OuyangMilzarek2024], the convergence analysis of a trust-region globalization of a semismooth Newton method applied to the normal map equation is developed in finite dimensions, where globalization is based on the residual in the normal equation.
@@
In [Graser2008] a globalization of the active-set strategy [BergouniouxItoKunisch1999] to solve discretized box constrained optimal control problems was proposed. However, the analysis does not carry over to the infinite-dimensional setting, as certain operators are not invertible in infinite dimensions [Graser2008, Remark 3].
@@
where $F:Y\to\mathbb{R}$ is convex. Let $F^{*}$ denote the convex conjugate of $F$. Under the assumption that $F$ is strongly convex and twice continuously differentiable, we get that $F^{*}$ is also strongly convex and twice continuously differentiable. Let us denote with $\nabla F^{*}:=\nabla(F^{*})$ its gradient. Then, we have $\nabla F(\nabla F^{*}(y))=y$ for all $y\in Y$ by [BauschkeCombettes2017, Theorem 16.29]. Differentiating this identity yields $\nabla^{2}F(\nabla F^{*}(y))\nabla^{2}F^{*}(y))=\operatorname{id}$ or $\nabla^{2}F^{*}(y)=\nabla^{2}F(\nabla F^{*}(y))^{-1}$. The resulting second-order derivative of $\Phi$ is given by $\partial^{2}\Phi=\nabla^{2}F^{*}+\partial^{2}\Psi$. Then all the convergence analysis for the corresponding semismooth Newton algorithm remains true, only the constants in Lemma˜5.7 and Corollary˜5.4 have to be adapted.
@@ 6 Semismoothness of Nemyzki operators
(6.1) $g(u)=\int_{\Omega}\tilde{g}(u(x))\,\text{{}{}d}\mu$
@@
Proof 6.2.
This follows from the results of [Rockafellar1976, Section 3], in particular the characterization of $\partial g(u)$ as
@@ 6.1 Piecewise smooth proximals
Let $\operatorname{prox}_{\tilde{g}/\alpha}$ satisfy ˜2. Define the single-valued map
@@
Proof 6.4.
Take $v\in\mathbb{R}$, and $h$ such that $\operatorname{prox}_{\tilde{g}/\alpha}$ is continuously differentiable between $v$ and $v+h$. Then both claims follow easily.
Clearly, this result can be generalized for $\operatorname{prox}_{\tilde{g}/\alpha}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with infinitely many pieces, see also [Ulbrich2011, Section 2.5.3].
In the next step, we prove that the derivative of the proximal operator of the integral functional $g$ is semismooth from $L^{q}(\Omega)$ to $L^{p}(\Omega)$ for $p<q$. The proof of the following result is adapted from the proof of [PotzlSchielaJaap2022, Proposition 6], more general results can be found in [Ulbrich2011, Section 3.3].
@@
Let $1\leq p<q\leq\infty$ be given. Let $\mu(\Omega)<\infty$. Let $\operatorname{prox}_{\tilde{g}/\alpha}$ satisfy ˜2. Let $v\in L^{q}(\Omega)$. Define the measurable function $\partial\operatorname{prox}_{g/\alpha}(v)$ for $\mu$-almost all $x\in\Omega$ by
@@
In addition, we set $r(x,0)=0$ for $x\in\Omega$. Then $x\mapsto r(x,h(x))$ is $\mu$-measurable for all measurable $h:\Omega\to\mathbb{R}$. In addition, $\lim_{h\to 0}r(x,h)=0$ for almost all $x$ by Lemma˜6.3. Since $\operatorname{prox}_{\tilde{g}/\alpha}$ is Lipschitz (Proposition˜2.5) and consequently $|\partial\operatorname{prox}_{\tilde{g}/\alpha}|\leq 1$, we have for almost all $x$ and all $h$ the estimate $|r(x,h)|\leq 2$.
@@
The second claim can be proven with similar arguments, observing that $|\operatorname{prox}_{\tilde{g}/\alpha}^{\prime}(v;h)|\leq|h|$.
Note that we cannot get rid of the norm gap in the previous result. The necessity of this gap has been discussed in [Ulbrich2011, Example 3.57], [ItoKunisch2008, Example 8.14] for semismoothness and in [GriesseGrundWachsmuth2008, Remark 4.4], [Malanowski2003] for Bouligand differentiability.
@@
Without the assumption $\mu(\Omega)<\infty$, the result of Lemma˜6.5 is not true. Let
@@
Due to the norm gap appearing in Lemma˜6.5, we cannot expect that $\operatorname{prox}_{g/\alpha}$ is semismooth as a mapping on $L^{2}(\Omega)$. In order that ˜1 is satisfied, we will need to rely on smoothing properties of $S$ or $S^{*}$.
@@
Assume $\mu(\Omega)<\infty$. Set $U:=L^{2}(\Omega)$. Let $\operatorname{prox}_{\tilde{g}/\alpha}$ satisfy ˜2. Let $Y$ be a Hilbert space, $S\in\mathcal{L}(L^{2}(\Omega),Y)$ be given such that the Hilbert space adjoint $S^{*}$ belongs to $\mathcal{L}(Y,L^{q}(\Omega))$ for some $q>2$.
@@
(6.2) $\partial^{2}\Psi(\xi):=\left\{\frac{1}{\alpha}S\partial\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)S^{*}\right\},$
where $\partial\operatorname{prox}_{g/\alpha}$ is defined in Lemma˜6.5. Moreover, ˜1 is satisfied.
@@
Proof 6.9 (Proof of Lemma˜6.8).
We show that $\partial^{2}\Psi(\xi)$ from (6.2) satisfies the assumptions of Lemma˜3.2. Let us recall $\nabla\Psi(\xi)=S\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha)$. Let $v,h\in Y$ be given. Then Lemma˜6.5 with $p=2$ and the assumption on $S^{*}$ imply
@@
Clearly this implies that $\partial^{2}\Psi(\xi)$ satisfies (3.2). The Bouligand property of the directional derivative can be proven with analogous arguments.